# Operator Learning for Continuous Spatial-Temporal Model with Gradient-Based and Derivative-Free Optimization Methods

### v2

@@ 2.5 Neural Dynamical Operator
Besides being good at predicting short-term system states, the trained modeled system in Eq. (2.2) also demonstrates a stable long-term simulation. There are some existing methods to make stable long-term predictions for chaotic dynamical systems such as reduced order modeling maulik2020time and stabilized neural ODE linot2023stabilized . Instead of preventing the high-wavenumbers amplification by dimension reduction or employing a linear damping term, the use of the Fourier neural operator in this work facilitates the stable long-term simulation by filtering out the high-order modes in the Fourier space. Therefore, neural dynamical operator trained via short-term time series data can still avoid numerical explosion in long-term simulations and also qualitatively retain the statistical property of the true system.
@@
Figure 2.1: Schematic diagram of continuous spatial-temporal model by neural dynamical operator (based on Navier-Stokes Equation). The dynamics of the system for the current state are approximated by neural operator, then the future states are evaluated with an ODE solver along with time with a given initial state. The neural dynamical operator $\tilde{\mathcal{G}}$ is trained by minimizing the Loss with gradient-based optimization.
@@ 2.6 A Hybrid Optimization Scheme
To efficiently incorporate both short-term time series and long-term statistics of the true system states as the training data, we propose a hybrid optimization scheme that iteratively solves the optimization problems in Eqs. (2.3) and (2.4). Within each iteration, the parameters $\theta$ of the dynamical operator $\tilde{\mathcal{G}}$ in Eq. (2.2) are first updated via a gradient-based optimization method (e.g., backpropagation or adjoint sensitivity) that solves Eq. (2.3) with the short-term time series data of the true system states and then further adjusted via a derivative-free optimization method (e.g., ensemble Kalman inversion) to account for the long-term statistics data in Eq. (2.4).
The key merit of the hybrid optimization scheme is the efficient incorporation of both short-term time series and long-term time-averaged statistics data. A detailed algorithm of the hybrid optimization scheme that leverages two types of data is presented in A. With the use of both types of data, the trained dynamical operator is expected to have better generalization capability, which is confirmed by the numerical example of the Kuramoto–Sivashinsky equation in Section 3.3.
Figure 2.2: Schematic diagram of continuous spatial-temporal model by neural dynamical operator with hybrid optimization scheme (based on Navier-Stokes equation). To better generalize the model by utilizing both short-term and long-term data, the neural dynamical operator $\tilde{\mathcal{G}}$ is trained by the hybrid optimization scheme which will iteratively update parameters by gradient-based method (SGD) to minimize short-term states loss $L$ and by derivative-free method (EKI) to minimize long-term statistics loss $L_{l}$. The short-term system evolution in $[t_{0},t_{N}]$ corresponds to Fig. 2.1.
@@ 3.1 Viscous Burgers’ Equation
Burgers’ equation is a classical example of partial differential equation and has been widely used in many fields such as fluid mechanics burgers1948mathematical ; hopf1950partial ; bec2007burgers , nonlinear dynamics kardar1986dynamic ; weinan2000invariant and traffic flow helbing2001traffic ; nagatani2002physics . The governing equation of the viscous Burgers’ equation is:
@@
The simulation settings are $L_{x}=1,L_{t}=5,\nu=10^{-3},dx=1/1024,dt=0.005$. We perform 1000 simulations as training data and another 100 simulations as test data, and the initial conditions of these simulations are randomly sampled from a 1-D Gaussian random field $\mathcal{N}(0,625(-\Delta+25I)^{-2})$ with periodic boundary conditions, where $\Delta$ is a Laplace operator and $I$ is an identity matrix. This choice of Gaussian random field is the same as in the paper of FNO li2021fourier .
In the optimization problem in Eq. (2.3), we first simulate the modeled system for $L_{t}=5$ time units given an initial state $u_{t_{0}}$ to obtain the trajectory $\{u_{t_{n}}\}_{n=1}^{N}$ and then minimize the $\ell^{2}$-norm between the simulated system trajectory and the true one. During the training process, we select a small subset from 1000 training simulations as a mini-batch to perform the gradient descent optimization in each epoch.
@@
Figure 3.3 presents the energy spectrum of the solution $u$ for the true system and the model ones. The spectrum shows a slope $k^{-2}$, which agrees well with linot2023stabilized . The three trained models are tested with the same resolution ($dx=1/1024$, $dt=0.05$). It can be seen that the energy spectrum of the trained models have good agreement with each other, indicating that the resolution-invariance is achieved by the proposed neural dynamical operator. In addition, we can see that the trained models can capture the true spectrum up to a wavenumber of approximately $200$ to $300$. The main reason for the noticeable difference from the true system for larger wavenumbers is that all three initial conditions are not included in the training data. It is promising to see that the trained models still generalize quite well with all these unseen initial conditions in most wavenumbers and only start to display mismatches for very high wavenumbers.
@@ 3.2 Navier–Stokes Equations
We consider the Navier–Stokes equations to study the performance of neural dynamical operator on a 2-D continuous dynamical system. The Navier-Stokes equations are partial differential equations that characterize the conservation of linear momentum in fluid flows navier1822memoire ; chorin1968numerical ; acheson1991elementary ; temam2001navier . The 2-D Navier–Stokes equations written in the form of vorticity are:
@@
Figure 3.4 presents the energy spectrum of data of initial condition from 1-D viscous Burgers’ equation and 2-D Navier-Stokes equation with respect to different resolution settings. We can see in Fig. 3.4(a) that the energy spectrum of the viscous Burgers’ equation is similar for most of the wave numbers with different spatial resolutions. However, Fig. 3.4(b) shows more noticeable differences in the energy spectrum of Navier–Stokes equations across the whole range of wave numbers with respect to different spatial resolutions. Unlike spatial discretization of VBE that all the resolution settings provide a consistent result of energy spectrum, the coarse resolution settings of NSE can cause over-estimations of energy in low wave numbers, which indicates that the numerical simulations do not well capture the true dynamical operator. Therefore, the trained models in the VBE example can approximate the true continuous operator and adapt well to the different resolution settings in Table. 3.1, with small test errors when making predictions in a higher spatial resolution. However, in the example of NSE, the trained models from coarse resolution settings in Table 3.2 provide large test errors when making predictions in a higher spatial resolution, mainly due to the information loss in high wave numbers that prevent a good approximation of the true continuous operator.
Figure 3.4: Energy spectrum of initial condition data of viscous Burgers’ equation and Navier-Stokes equation with respect to different resolution settings in the Table 3.1 and Table 3.2.
@@
We further compare the energy spectrum of the solutions between the true system and the trained models in Fig. 3.7. A reference slope $k^{-3}$ is also included, which corresponds to the empirical decay rate of 2-D turbulence based on experimental data. The three trained models are tested with the same spatial-temporal resolution (i.e., $dx=1/64,dy=1/64,dt=0.2$). We can see that only the energy spectrum of Model 3 has a good agreement with the true spectrum, while the results of Model 4 and 5 both demonstrate noticeable differences from the true one. It should be noted that the energy spectrum of the prediction results from Model 4 and 5 also do not agree well with the true results in Fig. 3.4(b), which indicates that the trained dynamical operator only based on data in very coarse resolutions may not generalize well to the finer resolutions.
Figure 3.7: Energy spectrum of 2-D Navier-Stocks equation simulated with an initial condition from test data at various times.
Figure 3.7: Energy spectrum of 2-D Navier-Stocks equation simulated with an initial condition from test data at various times.
@@ 3.3 Kuramoto–Sivashinsky Equation
Kuramoto–Sivashinsky (K-S) equation kuramoto1976persistent ; ashinsky1988nonlinear ; hyman1986kuramoto ; kevrekidis1990back is a fourth-order nonlinear partial differential equation that was originally developed to model diffusive–thermal instabilities in a laminar flame front and features chaotic behavior and rich dynamics, e.g., dissipation and dispersion. The governing equation of the K-S equation is:
@@
where $x\in(0,L_{x})$ with periodic boundary conditions, $t\in[0,L_{t}]$ and $u(x,0)$ is the given initial condition. We aim to learn a neural dynamical operator to approximate the right-hand-side of Eq. (3.4), i.e., $\mathcal{G}:u\mapsto\frac{\partial u}{\partial t}$. In this example, we demonstrate that (i) the trained neural dynamical operator is spatial-temporal resolution-invariant and can provide good short-term prediction, (ii) the trained model based on short-term data can provide a stable long-term simulation with the chaotic behavior being retained qualitatively, and (iii) the model can achieve good performance for both short-term trajectory prediction and long-term statistics matching when trained with a hybrid optimization method.
The simulation settings are $L_{x}=22,L_{t}=5000,dx=22/1024,dt=0.025$ and the initial condition is $u(x,0)=0.1\times\cos(x/16)\times(1+2\sin(x/16))$. We simulate the true system with a single long trajectory, and the first $80\%$ of the trajectory (4000 time units) is used as train data and the remaining $20\%$ (1000 time units) is used as test data.
@@
which is estimated from samples of both distributions based on k-Nearest-Neighbours probability density estimation wang2009divergence . In this work, the KL divergence is estimated from true states to predicted states for 1000 time units in test data. Each point in Fig. 3.12 stands for $D_{\textrm{KL}}$ from PDF of true data and PDF of predicted data.
We summarize the short-term test errors and long-term KL divergence for the system state $u$ in Table 3.3. The Test Error (I) and Long-Term $D_{\textrm{KL}}$ (I) are based on the test data of same resolution $dx=1/1024$ and $dt=0.05$, and the Test Error (II) and Long-Term $D_{\textrm{KL}}$ (II) is based on a resolution setting the same as the one of each train data. By comparing these test errors for trained models on different resolutions, it can be seen that the test error stays at the same order of magnitude when testing on a finer resolution in both space and time, confirming the resolution-invariance property of the trained models. In addition, all trained models lead to stable long-term simulations, which is mainly due to the high wavenumber filtering at each time step of evaluating the neural dynamical operator. More importantly, the stable long-term simulations of the trained models demonstrate small errors in KL divergence of the system state $u$, indicating a good quantitative agreement for the long-term prediction of the system state with the initial conditions from the test data. It should be noted that the long-term simulation results could still be inaccurate if the models are solely trained with short-term trajectory data. For instance, the relatively large Long-Term $D_{\textrm{KL}}$ (II) in the bottom row of Table 3.3 indicates a less satisfactory long-term prediction performance.
@@
Figure 3.8: Solution profiles of the K-S equation for the true system and the model ones with different initial conditions from test data. The model is trained in a coarse resolution ($dx=1/64,dt=0.5$) and tested on a finer resolution ($dx=1/1024,dt=0.05$).
@@
We further examine the statistical properties of long-term solutions from true and model systems in Fig. 3.10. Based on the true simulations and predictions for $1000$ time units, we compare the probability density function (PDF) of the system state $u$, first-order spatial derivative $u_{x}$, second-order spatial derivative $u_{xx}$, and spatial and temporal auto-correlation function (ACF) of $u$. We can find that the PDF of state $u$ and its first spatial derivative from long-term prediction can match the true simulation well, while the PDF of the second spatial derivative $u_{xx}$ from the modeled systems show a less satisfactory agreement with the true one. Although the mean and variance of $u_{xx}$ from each modeled system are still close to the true one, however, the PDFs of the modeled systems are less peaked around the mean, which indicates a lower kurtosis. Also, the temporal and spatial ACF of state $u$ show a similar pattern between predicted and true values, indicating a stable long-term prediction by trained models with similar statistical properties. In Fig. 3.11, we also show the joint probability density function of $(u_{x},u_{xx})$ for the long-term (1000 time units) simulation from true system and modeled systems with the same resolution $dx=22/1024,dt=0.25$. Compared with the joint PDF from true simulation, even though the joint PDF from model predictions have a relatively lower maximum density and spread more out, their overall pattern are still qualitatively similar to the true system.

### v3

@@ 2.5 Neural Dynamical Operator
Besides being good at predicting short-term system states, the trained modeled system in Eq. (2.2) also demonstrates a stable long-term simulation. There are some existing methods to make stable long-term predictions for chaotic dynamical systems such as reduced order modeling [48] and stabilized neural ODE [52]. Instead of preventing the high-wavenumbers amplification by dimension reduction or employing a linear damping term, the use of the Fourier neural operator in this work facilitates the stable long-term simulation by filtering out the high-order modes in the Fourier space. Therefore, neural dynamical operator trained via short-term time series data can still avoid a numerical explosion in long-term simulations and also qualitatively retain the statistical property of the true system.
@@
Figure 2.1: Schematic diagram of neural dynamical operator (based on Navier–Stokes equations). The dynamics of the system for the current state are approximated by neural operator, then the future states are evaluated with an ODE solver along with time given any initial state. The neural dynamical operator $\tilde{\mathcal{G}}$ is trained by minimizing the Loss $L_{s}$ with gradient-based optimization.
@@ 2.6 A Hybrid Optimization Scheme
To efficiently incorporate both short-term time series and long-term statistics of the true system states as the training data, we propose a hybrid optimization scheme that iteratively solves the optimization problems in Eqs. (2.3) and (2.4). Within each iteration, the parameters $\bm{\theta}$ of the dynamical operator $\tilde{\mathcal{G}}$ in Eq. (2.2) are first updated via a gradient-based optimization method (e.g., backpropagation) that solves Eq. (2.3) with the short-term time series data of the true system states and then further adjusted via a derivative-free optimization method (e.g., ensemble Kalman inversion) to account for the long-term statistics data in Eq. (2.4). The Algorithm 1 briefly summarizes the hybrid optimization scheme, while a detailed version is presented in Algorithm 2 in A.
Algorithm 1 Training neural dynamical operator with the hybrid optimization scheme (A brief version)
1:for $i\leftarrow 1$ to $N_{\textrm{ep}}$ do
2:     $\bm{\theta}\leftarrow\bm{\theta}-\lambda\cdot\nabla_{\bm{\theta}}L_{s}$triangleright SGD Updating
3:     if  $i\bmod k=0$  then
4:         for $m\leftarrow 1$ to $N_{\textrm{it}}$ do
5:              $\{\bm{\theta}^{(j)}\}_{j=1}^{J}\leftarrow\{\bm{\theta}^{(j)}+\bm{\Sigma}^{\bm{\theta}\bm{g}}(\bm{\Sigma}^{\bm{gg}}+\bm{\Sigma}_{\bm{\eta}})^{-1}(\bm{y}^{(j)}-\bm{g}^{(j)})\}_{j=1}^{J}$triangleright EKI Updating
6:         end for
7:         $\bm{\theta}\leftarrow\frac{1}{J}\Sigma_{j=1}^{J}\bm{\theta}^{(j)}$
8:     end if
9:end for
The key merit of the hybrid optimization scheme is the efficient incorporation of both short-term time series and long-term time-averaged statistics data. With the use of both types of data, the trained dynamical operator is expected to have better generalization capability, which is confirmed by the numerical example of the Kuramoto–Sivashinsky equation in Section 3.3. One thing should be noted is that the covariance matrix $\bm{\Sigma}_{\bm{\eta}}$ in Eq. (2.11) of EKI for long-term statistics matching does not affect the training for short-term trajectory matching, since the two optimization processes are independent of each other as shown in Algorithm 1. A simple choice for $\bm{\Sigma}_{\bm{\eta}}$ is a diagonal matrix, e.g., $c^{2}\bm{\mathrm{I}}_{d_{\bm{y}}}$ where $c\in\mathbb{R}$ is a constant scalar and $\bm{\mathrm{I}}_{d_{\bm{y}}}\in\mathbb{R}^{d_{\bm{y}}\times d_{\bm{y}}}$ is an identity matrix with dimension $d_{\bm{y}}$.
Figure 2.2: Schematic diagram of neural dynamical operator with hybrid optimization scheme (based on Navier–Stokes equations). To better generalize the model by utilizing both short-term and long-term data, the neural dynamical operator $\tilde{\mathcal{G}}$ is trained by the hybrid optimization scheme which will iteratively update parameters by stochastic gradient descent (SGD) method to minimize short-term states loss $L_{s}$ and by derivative-free method (EKI) to minimize long-term statistics loss $L_{l}$. The short-term system evolution in $[t_{0},t_{N_{s}}]$ corresponds to Fig. 2.1.
@@ 3.1 Viscous Burgers’ Equation
Burgers’ equation is a classical example of partial differential equation and has been widely used in many fields such as fluid mechanics [74, 75, 76], nonlinear dynamics [77, 78] and traffic flow [79, 80]. The governing equation of the viscous Burgers’ equation is:
@@
The simulation settings are $L_{x}=1,L_{t}=5,\nu=10^{-3},dx=1/1024,dt=0.005$. We perform 1000 simulations as training data and another 100 simulations as test data, and the initial conditions of these simulations are randomly sampled from a 1-D Gaussian random field $\mathcal{N}(0,625(-\Delta+25I)^{-2})$ with periodic boundary conditions, where $\Delta$ is a Laplace operator and $I$ is an identity operator. This choice of Gaussian random field is the same as in the paper of FNO [28].
In the optimization problem in Eq. (2.3), we first simulate the modeled system for $L_{t}=5$ time units given an initial state $u(t_{0})$ to obtain the trajectory $\{\tilde{u}(t_{n})\}_{n=1}^{N_{s}}$ with $t_{N_{s}}=5$ and then minimize the $\ell^{2}$-norm between the simulated system trajectory and the true one. During the training process, we select a small subset from 1000 training simulations as a mini-batch to perform the gradient descent optimization in each epoch.
@@
Figure 3.3 presents the energy spectrum of the solution $u$ for the true system and the model ones. The spectrum shows a slope $k^{-2}$, which agrees well with [52]. The three trained models are tested with the same resolution ($dx=1/1024$, $dt=0.05$). It can be seen that the energy spectrum of the trained models has good agreement with each other, indicating that the resolution-invariance is achieved by the proposed neural dynamical operator. In addition, we can see that the trained models can capture the true spectrum up to a wavenumber of approximately $200$ to $300$. The main reason for the noticeable difference from the true system for larger wavenumbers is that all three initial conditions are not included in the training data. It is promising to see that the trained models still generalize quite well with all these unseen initial conditions in most wavenumbers and only start to display mismatches for very high wavenumbers.
@@ 3.2 Navier–Stokes Equations
We consider the Navier–Stokes equations to study the performance of neural dynamical operator on a 2-D continuous dynamical system. The Navier–Stokes equations are partial differential equations that characterize the conservation of linear momentum in fluid flows [81, 82, 83, 84]. The 2-D Navier–Stokes equations written in the form of vorticity are:
@@
Figure 3.4 presents the energy spectrum of data of initial condition from 1-D viscous Burgers’ equation and 2-D Navier–Stokes equation with respect to different resolution settings. We can see in Fig. 3.4(a) that the energy spectrum of the viscous Burgers’ equation is similar for most of the wave numbers with different spatial resolutions. However, Fig. 3.4(b) shows more noticeable differences in the energy spectrum of Navier–Stokes equations across the whole range of wave numbers with respect to different spatial resolutions. Unlike spatial discretization of VBE, where all the resolution settings provide a consistent result of the energy spectrum, the coarse resolution settings of NSE can cause over-estimations of energy in low wave numbers, which indicates that the numerical simulations do not well capture the true dynamical operator. Therefore, the trained models in the VBE example can approximate the true continuous operator and adapt well to the different resolution settings in Table. 3.1, with small test errors when making predictions in a higher spatial resolution. However, in the example of NSE, the trained models from coarse resolution settings in Table 3.2 provide large test errors when making predictions in a higher spatial resolution, mainly due to the information loss in high wave numbers that prevent a good approximation of the true continuous operator.
Figure 3.4: Energy spectrum of initial condition data of viscous Burgers’ equation and Navier–Stokes equation with respect to different resolution settings in the Table 3.1 and Table 3.2.
@@
We further compare the energy spectrum of the solutions between the true system and the trained models in Fig. 3.7. A reference slope $k^{-3}$ is also included, which corresponds to the empirical decay rate of 2-D turbulence based on experimental data. The three trained models are tested with the same spatial-temporal resolution (i.e., $dx=1/64,dy=1/64,dt=0.2$). We can see that only the energy spectrum of Model 3 has a good agreement with the true spectrum, while the results of Models 4 and 5 both demonstrate noticeable differences from the true one. It should be noted that the energy spectrum of the prediction results from Models 4 and 5 also do not agree well with the true results in Fig. 3.4(b), which indicates that the trained dynamical operator only based on data in very coarse resolutions may not generalize well to the finer resolutions.
Figure 3.7: Energy spectrum of 2-D Navier–Stokes equation simulated with an initial condition from test data at various times.
@@ 3.3 Kuramoto–Sivashinsky Equation
Kuramoto–Sivashinsky (K-S) equation [85, 86, 87, 88] is a fourth-order nonlinear partial differential equation that was originally developed to model diffusive–thermal instabilities in a laminar flame front and features chaotic behavior and rich dynamics, e.g., dissipation and dispersion. The governing equation of the K-S equation is:
@@
where $x\in(0,L_{x})$ with periodic boundary conditions, $t\in[0,L_{t}]$ and $u(x,0)$ is the given initial condition. We aim to learn a neural dynamical operator to approximate the right-hand-side of Eq. (3.4), i.e., $\mathcal{G}:u\mapsto\frac{\partial u}{\partial t}$. In this example, we demonstrate that (i) the trained neural dynamical operator is spatial-temporal resolution-invariant and can provide good short-term prediction, (ii) the trained model based on short-term data can provide a stable long-term simulation with the chaotic behavior being retained qualitatively, and (iii) the model can achieve good performance for both short-term trajectory prediction and long-term statistics matching when trained with a hybrid optimization method. These three features are described in detail in the subsequent sections.
The simulation settings are $L_{x}=22,L_{t}=5000,dx=22/1024,dt=0.025$ and the initial condition is $u(x,0)=0.1\times\cos(x/16)\times(1+2\sin(x/16))$. We simulate the true system with a single long trajectory, and the first $80\%$ of the trajectory (4000 time units) is used as train data and the remaining $20\%$ (1000 time units) is used as test data. For the K-S equation, 1000 time units is long enough to demonstrate the stability of long-term prediction of neural dynamical operator.
3.3.1 Short-Term Prediction with Spatial-Temporal Resolution-Invariant
@@
which is estimated from samples of both distributions based on k-Nearest-Neighbours probability density estimation [89]. In this work, the KL divergence is estimated from true states to predicted states for 1000 time units in test data. Each point in Fig. 3.12 stands for $D_{\textrm{KL}}$ from PDF of true data and PDF of predicted data.
We summarize the short-term test errors and long-term KL divergence for the system state $u$ in Table 3.3. The Test Error (I) and Long-Term $D_{\textrm{KL}}$ (I) are based on the test data of same resolution $dx=1/1024$ and $dt=0.05$, and the Test Error (II) and Long-Term $D_{\textrm{KL}}$ (II) is based on a resolution setting the same as the one of each train data. By comparing these test errors for trained models on different resolutions, it can be seen that the test error stays at the same order of magnitude when testing on a finer resolution in both space and time, confirming the resolution-invariance property of the trained models. In addition, all trained models lead to stable long-term simulations, which is mainly due to the high wavenumber filtering at each time step of evaluating the neural dynamical operator. More importantly, the stable long-term simulations of the trained models demonstrate small errors in KL divergence of the system state $u$, indicating a good quantitative agreement for the long-term prediction of the system state with the initial conditions from the test data. It should be noted that the long-term simulation results could still be inaccurate if the models are solely trained with short-term trajectory data. For instance, the relatively large Long-Term $D_{\textrm{KL}}$ (II) in the bottom row of Table 3.3 indicates a less satisfactory long-term prediction performance. In addition, the lowest Long-Term $D_{\textrm{KL}}$ (I) from model trained with Resolution 3 in Table 3.3 is mainly due to the resolution-invariant feature of the neural dynamical operator. The visualization of $D_{\textrm{KL}}$ (I) is shown in the first plot of Figure 3.10. It can be observed that the probability density functions (PDFs) obtained from models trained with the three different resolutions are very similar.
@@
Figure 3.8: Solution profiles of the K-S equation for the true system and the model ones with different initial conditions from test data. The model is trained with each resolutions in Table. 3.3 and tested on a finer resolution ($dx=22/1024,dt=0.25$).
3.3.2 Stable Long-Term Simulation with Chaotic Behavior
@@
We further examine the statistical properties of long-term solutions from true and model systems in Fig. 3.10. Based on the true simulations and predictions for $1000$ time units, we compare the probability density function (PDF) of the system state $u$, first-order spatial derivative $u_{x}$, second-order spatial derivative $u_{xx}$, and spatial and temporal auto-correlation function (ACF) of $u$. We can find that the PDF of state $u$ and its first spatial derivative from long-term prediction can match the true simulation well, while the PDF of the second spatial derivative $u_{xx}$ from the modeled systems shows a less satisfactory agreement with the true one. Although the mean and variance of $u_{xx}$ from each modeled system are still close to the true one, however, the PDFs of the modeled systems are less peaked around the mean, which indicates a lower kurtosis. Also, the temporal and spatial ACF of state $u$ show a similar pattern between predicted and true values, indicating a stable long-term prediction by trained models with similar statistical properties. In Fig. 3.11, we also show the joint probability density function of $(u_{x},u_{xx})$ for the long-term (1000 time units) simulation from true system and modeled systems with the same resolution $dx=22/1024,dt=0.25$. Compared with the joint PDF from true simulation, even though the joint PDFs from model predictions have a relatively lower maximum density and spread more out, their overall patterns are still qualitatively similar to the true system.
@@
3.3.3 Hybrid Optimization for Short-Term State Prediction and Long-Term Statistics Matching